### v1

@@ 1 Introduction
Text-to-image (T2I) diffusion models, such as Stable Diffusion (SD) (Rombach et al., 2022; Podell et al., 2023) and DALL-E (Ramesh et al., 2022), have garnered substantial attention for their ability to synthesize high-quality images from natural language prompts through iterative denoising and cross-modal attention mechanisms. These models have been adopted in a wide range of applications, including image editing (Huang et al., 2024b; Kawar et al., 2023; Liu et al., 2024a; Mou et al., 2024), data augmentation (Li et al., 2024c; Xiao et al., 2023; Feng et al., 2023a), medical imaging (Huang et al., 2024a; Li et al., 2024b; Khader et al., 2023; Lin et al., 2024a), and marketing (Shilova et al., 2023; Yang et al., 2024). Despite their versatility and impressive generation capabilities, T2I diffusion models often exhibit notable failures in compositional alignment (Huang et al., 2025; Bakr et al., 2023; Ghosh et al., 2024). These failures manifest in various forms, including entity omission (Chefer et al., 2023b; Sueyoshi & Matsubara, 2024a; Zhang et al., 2024a; Liu et al., 2022; Kim et al., 2023), incorrect attribute binding (Feng et al., 2023b; Li et al., 2023c; Rassin et al., 2024; Wang et al., 2024), misrepresentation of spatial relationships (Chatterjee et al., 2024; Gokhale et al., 2022; Chen et al., 2024), and numeracy errors (Binyamin et al., 2024; Zafar et al., 2024; Kang et al., 2023b).
@@ 2 Related Works
Research on compositional generation in T2I diffusion models can be grouped into two families: fine-tuning methods, which update model parameters, and inference-time methods, which enhance alignment without additional training. Fine-tuning either modifies the denoiser or the text encoder. Denoiser-level updates adapt the UNet or add auxiliary modules for spatial control and attribute binding Sun et al. (2023); Jiang et al. (2024); Guo et al. (2024a); Zhang et al. (2023); Mou et al. (2023), but demand extra compute and risk overfitting. Encoder-level fine-tuning instead adjusts the conditioning space with lightweight projections or causal refinements over frozen CLIP embeddings Zarei et al. (2024; 2025), offering better generalization but weaker handling of spatial errors. Inference-time methods operate at different stages of generation. Prompt-level rewriting with lexical search or LLM feedback improves attributes and personalization Yu et al. (2024); He et al. (2025), though often costly and verbose. Embedding-level adjustments refine frozen encoders to control object counts, attributes, or relations Smith et al. (2024); Deckers et al. (2024), but are sensitive to hyperparameters. Finally, noise- and latent-level methods exploit the strong influence of initialization, forming the basis for optimization and exploration strategies detailed in the next subsections.
@@
From this category, ImageSelect  Karthik et al. (2023) and SeedSelect  Samuel et al. (2024b) search over candidate seeds, choosing the one best matching a scoring heuristic (e.g., CLIP similarity). SemI  Mao et al. (2024) biases selection toward noise vectors empirically linked to stronger object binding, exploiting “lucky” seeds as reproducible advantages. ParticleFiltering  Liu et al. (2024b) instead performs sequential resampling during reverse diffusion, discarding low-scoring partial generations and retaining promising ones. These methods are fully training-free and turn stochastic seed choice into systematic search, but incur high computational cost, depend on potentially noisy scoring signals, and remain insufficient when base models exhibit severe binding failures.
@@
These methods mainly instead refine the initial noise iteratively using reward signals from the final generated image, directly enforcing compositional constraints at test time. InitNO Guo et al. (2024b) optimizes noise with an attention-aware objective that penalizes missing objects and concept mixing, steering sampling away from neglect-inducing regions. ReNO Eyring et al. (2024) extends this to multi-reward optimization, ascending gradients of preference models (e.g., text alignment or detection scores) with respect to noise, improving counting, co-occurrence, and attribute binding. These methods provide strong, training-free gains and flexibly integrate new constraints, but add inference overhead from iterative scoring/backpropagation, remain sensitive to reward design (risk of reward hacking), and rely on external scorer quality.
@@
The category of methods refine noisy latent codes using loss functions on intermediate cross-attention maps, encouraging concept preservation and disentanglement during denoising. Attend-and-Excite Chefer et al. (2023b) mitigates neglect by amplifying subject-token activations, while Divide&Bind Li et al. (2023d) adds attendance and binding losses for multi-entity prompts and attribute–object pairing. Predicated Diffusion Sueyoshi & Matsubara (2024b) encodes prompt semantics as predicate-logic propositions and treats attention maps as fuzzy predicates, enabling differentiable objectives for complex relations. Attention Regulation Zhang et al. (2024b) formulates cross-attention control as constrained optimization that suppresses dominant tokens and boosts under-attended ones, and A-STAR Agarwal et al. (2023) combines attention segregation (reducing token overlap) with retention (preserving salience across timesteps). Collectively, these training-free methods improve semantic fidelity, recall, and binding by targeting the attention interface, but add inference overhead and are sensitive to hyperparameters balancing faithfulness, diversity, and runtime.
@@ 2.1 Reward Models for Text-Image Alignment
Compositional alignment rewards assess how well a generated image matches a text prompt in terms of objects, attributes, and spatial relations. Embedding-based methods are widely used, with CLIPScore computing similarity between CLIP embeddings of text and image Hessel et al. (2021). Extensions include HPS, which fine-tunes CLIP on preference data to better match human judgments Wu et al. (2023), and PickScore, which adapts CLIP-H with preference supervision for closer correlation with human rankings Kirstain et al. (2023). Moreover, ImageReward trains a standalone reward model on human evaluations to capture prompt relevance and perceptual quality Xu et al. (2024). Complementary to these, VQA-based methods assess alignment by testing whether the image supports answers to prompt-derived questions: TIFA generates structured QA pairs and checks them with a pretrained VQA model Hu et al. (2023), while VQAScore applies a similar principle and achieves higher correlation with human judgments Lin et al. (2024b). Other approaches in this family, such as DA, DSG, and B-VQA Singh & Zheng (2023); Cho et al. (2023); Huang et al. (2023), also rely on question–answer correctness to provide compositional faithfulness scores.
@@ 3 Analysis of Optimization and Exploration Limitations
4 CARINOX: Reward-Guided Noise Optimization and Exploration
@@ 4.1.1 Gradient-Based Initial Noise Optimization
where $\lambda_{i}$ are the fixed weights assigned to each reward function, set to $1$ for all rewards in our implementation. This procedure ensures that no single reward term overwhelms the optimization, allowing for stable and balanced gradient-based updates (see Section E for experimental analysis).
@@ 4.2 Correlation-Guided Reward Combination Selection
To guide this choice, we conducted a systematic correlation study on the T2I-CompBench++ dataset (Huang et al., 2023), which provides curated prompts, generated images, and human evaluation scores. We tested five embedding-based metrics (PickScore (Kirstain et al., 2023), CLIPScore (Hessel et al., 2021), HPS (Wu et al., 2023), ImageReward (Xu et al., 2024), BLIP-2 (Li et al., 2023a)), five VQA-based metrics (B-VQA (Huang et al., 2023), DA Score (Singh & Zheng, 2023), TIFA (Hu et al., 2023), DSG (Cho et al., 2023), VQA Score (Lin et al., 2024b)), and two image-only metrics (CLIP-IQA (Wang et al., 2023), Aesthetic Score (Schuhmann et al., 2022)). Spearman rank correlation was used to assess alignment with human preference annotations across compositional categories (Table 5).
The study yielded several insights. No single metric was consistently optimal: performance varied across attributes, spatial relations, and numeracy. CLIPScore, despite widespread use, never ranked among the top metrics, underscoring its weakness as a standalone reward. VQA-based metrics showed strong compositional reasoning but were not uniformly superior across categories. Embedding-based metrics such as HPS and ImageReward frequently appeared among the top performers, showing correlations comparable to VQA-based scores. As expected, image-only metrics correlated poorly with human judgments.
@@
5 Experiments & Results
5.1 Experimental Setup
We evaluate CARINOX through a series of experiments designed to assess both compositional alignment and broader generation quality. Human evaluation (Section 5.2) provides direct judgments of alignment quality across multiple backbones. Automated evaluation on T2I-CompBench++ (Section 5.3) measures performance across diverse compositional categories, while the HRS benchmark (Section 5.4) extends this analysis to higher-level aspects such as creativity, style, and visual writing. For clarity, we compare three variants of our approach: CARINX, which applies our fixed reward combination for best-of-$N$ exploration; CARINO, which performs initial noise optimization with our reward-guided pipeline; and CARINOX, the full method that integrates both exploration and optimization. Together, these benchmarks and variants provide a comprehensive view of how our framework compares to baselines and state-of-the-art methods.
5.2 Human-Centered Assessment of Text–Image Alignment
We conducted a human study to directly assess the effectiveness of our proposed variants in improving text–image alignment. A set of 200 prompts covering all eight compositional categories of T2I-CompBench++ was used with two backbones, SD-Turbo and SDXL-Turbo. For each prompt, human annotators rated the generated images on a 0–3 scale, where scores reflected whether all objects were present, attributes were correctly rendered, and relations such as size, color, numeracy, and spatial layout were faithfully captured. Scores were averaged across raters, normalized to $[0,1]$, and reported in Table 1. Full details of the protocol are provided in Appendix F.
@@ 5.2 Human-Centered Assessment of Text–Image Alignment
5.3 Category-Level Compositional Benchmarking — T2I-CompBench++
We further benchmark our approach against a broad set of state-of-the-art methods on T2I-CompBench++, which evaluates eight compositional categories using specialized evaluators for each dimension. Since different metrics are used across categories (e.g., CLIP for non-spatial relations), the absolute ranges vary, but the comparisons remain consistent across methods. The baseline set includes standard diffusion backbones (SD v1.4, SD v2.1, SDXL, PixArt-$\alpha$), recent commercial models such as DALL·E 3, attention-based inference methods like Structured Diffusion Feng et al. (2022) and Attend-and-Excite Chefer et al. (2023b), noise optimization methods such as InitNO Guo et al. (2024b) and ReNO Eyring et al. (2024), and exploration-based approaches including Pick-a-Pic Kirstain et al. (2023) and ImageSelect Xu et al. (2024). We report results for three variants of our method: CARINX (exploration only), CARINO (optimization only), and CARINOX (combined).
Importantly, CARINO and CARINX also provide consistent improvements when used independently: CARINX surpasses existing exploration methods, while CARINO establishes a new strong baseline for noise optimization. Together, the full CARINOX pipeline further amplifies these gains, outperforming commercial systems such as DALL·E 3 and advancing the state of the art across compositional categories.
@@ 5.3 Category-Level Compositional Benchmarking — T2I-CompBench++
Table 2: Quantitative evaluation on T2I-CompBench++ across eight compositional categories using three different backbones. Results are reported for baseline models, state-of-the-art methods, and our variants. CARINOX achieves the strongest overall alignment, consistently surpassing both optimization- and exploration-based baselines. Best values are in bold, and second-best are underlined.
5.4 Beyond Compositionality: Expressive Evaluation on HRS
@@ 6.2 Evaluation of Individual Reward Functions
Table 4: Ablation of optimization iterations (a) and exploration seeds (b) on T2I-CompBench++; gains plateau after 50 iterations and 5 seeds, which we adopt as defaults.
@@ C.1 Evaluation Metrics
DA Score (Singh & Zheng, 2023) 0.772 0.463 0.711 0.318 0.453 0.488 0.297 0.462
@@
DA Score (Singh & Zheng, 2023) 0.603 0.337 0.534 0.247 0.357 0.364 0.206 0.347
@@
DA Score (Singh & Zheng, 2023) 4
@@
VQA-based metrics assess compositional alignment by casting text–image consistency as a question answering task. Questions derived from the prompt are posed to a pretrained VQA model, with scores based on the correctness of its responses. VQAScore (Lin et al., 2024b) generates yes/no questions from the text, while TIFA (Hu et al., 2023) uses structured templates to cover objects, attributes, and relations. Variants target specific aspects: DA Score (Singh & Zheng, 2023) asks entity–attribute questions to test binding, DSG (Cho et al., 2023) converts the text into a scene graph to verify entities and relations, and B-VQA (Huang et al., 2023) decomposes the text into object–attribute pairs, querying each with BLIP-VQA and combining the probabilities.

### v2

@@ 1 Introduction
Text-to-image (T2I) diffusion models, such as Stable Diffusion (SD) (Rombach et al., 2022; Podell et al., 2023) and DALL-E (Ramesh et al., 2022), have garnered substantial attention for their ability to synthesize high-quality images from natural language prompts through iterative denoising and cross-modal attention mechanisms. These models have been adopted in a wide range of applications, including image editing (Huang et al., 2024b; Kawar et al., 2023; Liu et al., 2024a; Mou et al., 2024), data augmentation (Li et al., 2024c; Xiao et al., 2023; Feng et al., 2023a), medical imaging (Huang et al., 2024a; Li et al., 2024b; Khader et al., 2023; Lin et al., 2024a), and marketing (Shilova et al., 2023; Yang et al., 2024). Despite their versatility and impressive generation capabilities, T2I diffusion models often exhibit notable failures in compositional alignment (Huang et al., 2025; Bakr et al., 2023; Ghosh et al., 2024). These failures manifest in various forms, including entity omission (Chefer et al., 2023; Sueyoshi and Matsubara, 2024; Zhang et al., 2024; Liu et al., 2022; Kim et al., 2023), incorrect attribute binding (Feng et al., 2023b; Singh and Zheng, 2023; Rassin et al., 2024; Wang et al., 2024), misrepresentation of spatial relationships (Zhang et al., 2024; Gokhale et al., 2022; Chen et al., 2024), and numeracy errors (Binyamin et al., 2024; Zafar et al., 2024; Kang et al., 2023b).
@@ 2 Related Works
Research on compositional generation in T2I diffusion models can be grouped into two families: fine-tuning methods, which update model parameters, and inference-time methods, which enhance alignment without additional training. Fine-tuning either modifies the denoiser or the text encoder. Denoiser-level updates adapt the UNet or add auxiliary modules for spatial control and attribute binding Sun et al. (2023); Jiang et al. (2024); Guo et al. (2024a); Zhang et al. (2023); Mou et al. (2023), but demand extra compute and risk overfitting. Encoder-level fine-tuning instead adjusts the conditioning space with lightweight projections or causal refinements over frozen CLIP embeddings Zarei et al. (2025), offering better generalization but weaker handling of spatial errors. Inference-time methods operate at different stages of generation. Prompt-level rewriting with lexical search or LLM feedback improves attributes and personalization Yu et al. (2024); He et al. (2025), though often costly and verbose. Embedding-level adjustments refine frozen encoders to control object counts, attributes, or relations Zafar et al. (2024); Deckers et al. (2024), but are sensitive to hyperparameters. Finally, noise- and latent-level methods exploit the strong influence of initialization, forming the basis for optimization and exploration strategies detailed in the next subsections.
2.1 Inference-Time Latent-Space Search and Optimization
@@
From this category, ImageSelect  Karthik et al. (2023) and SeedSelect  Samuel et al. (2024b) search over candidate seeds, choosing the one best matching a scoring heuristic (e.g., CLIP similarity). SemI  Mao et al. (2024) biases selection toward noise vectors empirically linked to stronger object binding, exploiting “lucky” seeds as reproducible advantages. ParticleFiltering  Liu et al. (2024b) instead performs sequential resampling during reverse diffusion, discarding low-scoring partial generations and retaining promising ones. Relatedly, recent inference-time scaling methods formulate reward alignment as Sequential Monte Carlo (SMC) sampling with a population of particles and sequential resampling: DAS Kim et al. (2025b) performs test-time alignment of diffusion models via SMC-based sampling to improve reward while mitigating over-optimization, Kim et al. (2025a) extend particle-based scaling to flow models by injecting stochasticity and reallocating compute via rollover budget forcing, and $\Psi$-Sampler Yoon et al. (2025) improves SMC-based alignment in score models by initializing particles from a reward-aware posterior rather than a Gaussian prior. These methods are fully training-free and turn stochastic seed choice into systematic search, but incur high computational cost, depend on potentially noisy scoring signals, and remain insufficient when base models exhibit severe binding failures.
@@
These methods mainly instead refine the initial noise iteratively using reward signals from the final generated image, directly enforcing compositional constraints at test time. InitNO Guo et al. (2024b) optimizes noise with an attention-aware objective that penalizes missing objects and concept mixing, steering sampling away from neglect-inducing regions. ReNO Eyring et al. (2024) extends this to multi-reward optimization, ascending gradients of preference models (e.g., text alignment or detection scores) with respect to noise, improving counting, co-occurrence, and attribute binding. Beyond these, DOODL Wallace et al. (2023) performs end-to-end optimization of the initial diffusion latent by backpropagating losses defined on the final generated image through the denoising process. Related source-latent optimization has also been explored for flow-based generators: D-Flow Ben-Hamu et al. (2024) differentiates through flow-based generation to optimize the source noise under arbitrary costs, and ORIGEN Min et al. (2025) applies reward-guided updates of the initial latent in a one-step flow model to enforce 3D orientation grounding. These methods provide strong, training-free gains and flexibly integrate new constraints, but add inference overhead from iterative scoring/backpropagation, remain sensitive to reward design (risk of reward hacking), and rely on external scorer quality.
@@
The category of methods refine noisy latent codes using loss functions on intermediate cross-attention maps, encouraging concept preservation and disentanglement during denoising. Attend-and-Excite Chefer et al. (2023) mitigates neglect by amplifying subject-token activations, while Divide&Bind Li et al. (2023c) adds attendance and binding losses for multi-entity prompts and attribute–object pairing. Predicated Diffusion Sueyoshi and Matsubara (2024) encodes prompt semantics as predicate-logic propositions and treats attention maps as fuzzy predicates, enabling differentiable objectives for complex relations. Attention Regulation Zhang et al. (2024) formulates cross-attention control as constrained optimization that suppresses dominant tokens and boosts under-attended ones, and A-STAR Agarwal et al. (2023) combines attention segregation (reducing token overlap) with retention (preserving salience across timesteps). Collectively, these training-free methods improve semantic fidelity, recall, and binding by targeting the attention interface, but add inference overhead and are sensitive to hyperparameters balancing faithfulness, diversity, and runtime.
@@ 2.2 Reward Models for Text-Image Alignment
Compositional alignment rewards assess how well a generated image matches a text prompt in terms of objects, attributes, and spatial relations. Embedding-based methods are widely used, with CLIPScore computing similarity between CLIP embeddings of text and image Hessel et al. (2021). Extensions include HPS, which fine-tunes CLIP on preference data to better match human judgments Wu et al. (2023), and PickScore, which adapts CLIP-H with preference supervision for closer correlation with human rankings Kirstain et al. (2023). Moreover, ImageReward trains a standalone reward model on human evaluations to capture prompt relevance and perceptual quality Xu et al. (2024). Complementary to these, VQA-based methods assess alignment by testing whether the image supports answers to prompt-derived questions: TIFA generates structured QA pairs and checks them with a pretrained VQA model Hu et al. (2023), while VQAScore applies a similar principle and achieves higher correlation with human judgments Lin et al. (2024b). Other approaches in this family, such as DA, DSG, and B-VQA Singh and Zheng (2023); Cho et al. (2023); Huang et al. (2023), also rely on question–answer correctness to provide compositional faithfulness scores.
@@ 3 Analysis of Optimization and Exploration Limitations
4 CARINOX: Reward-Guided Initial Noise Optimization and Exploration
@@ 4.1.1 Gradient-Based Initial Noise Optimization
where $\lambda_{i}$ are the fixed weights assigned to each reward function, set to $1$ for all rewards in our implementation. We additionally report an ablation on adaptive reward weighting in Appendix F. This procedure ensures that no single reward term overwhelms the optimization, allowing for stable and balanced gradient-based updates (see Appendix E for experimental analysis).
@@ 4.2 Correlation-Guided Reward Combination Selection
To guide this choice, we conducted a systematic correlation study on the T2I-CompBench++ dataset (Huang et al., 2023), which provides curated prompts, generated images, and human evaluation scores. We tested five embedding-based metrics (PickScore (Kirstain et al., 2023), CLIPScore (Hessel et al., 2021), HPS (Wu et al., 2023), ImageReward (Xu et al., 2024), BLIP-2 (Li et al., 2023a)), five VQA-based metrics (B-VQA (Huang et al., 2023), DA Score (Singh and Zheng, 2023), TIFA (Hu et al., 2023), DSG (Cho et al., 2023), VQA Score (Lin et al., 2024b)), and two image-only metrics (CLIP-IQA (Wang et al., 2023), Aesthetic Score (Schuhmann et al., 2022)). Spearman rank correlation was used to assess alignment with human preference annotations across compositional categories (Table 5).
The study yielded several insights Kasaei et al. (2025). No single metric was consistently optimal: performance varied across attributes, spatial relations, and numeracy. CLIPScore, despite widespread use, never ranked among the top metrics, underscoring its weakness as a standalone reward. VQA-based metrics showed strong compositional reasoning but were not uniformly superior across categories. Embedding-based metrics such as HPS and ImageReward frequently appeared among the top performers, showing correlations comparable to VQA-based scores. As expected, image-only metrics correlated poorly with human judgments.
@@
5 Experiments & Results
5.1 Experimental Setup
We evaluate CARINOX through a series of experiments designed to assess both compositional alignment and broader generation quality. Human evaluation (Section 5.2) provides direct judgments of alignment quality across multiple backbones. Automated evaluation on T2I-CompBench++ (Section 5.3) measures performance across diverse compositional categories, while the HRS benchmark (Section 5.4) extends this analysis to higher-level aspects such as creativity, style, and visual writing.
For clarity, we compare three variants of our approach: CARINX, which applies our fixed reward combination for best-of-$N$ exploration; CARINO, which performs initial noise optimization with our reward-guided pipeline; and CARINOX, the full method that integrates both exploration and optimization. Together, these benchmarks and variants provide a comprehensive view of how our framework compares to baselines and state-of-the-art (SOTA) methods.
5.2 Human-Centered Assessment of Text–Image Alignment
We conducted a human study to directly assess the effectiveness of our proposed variants in improving text–image alignment. A set of 200 prompts covering all eight compositional categories of T2I-CompBench++ was used with two backbones, SD-Turbo and SDXL-Turbo. For each prompt, human annotators rated the generated images on a 0–3 scale, where scores reflected whether all objects were present, attributes were correctly rendered, and relations such as size, color, numeracy, and spatial layout were faithfully captured. Scores were averaged across raters, normalized to $[0,1]$, and reported in Table 1. Full details of the protocol are provided in Appendix G.
@@ 5.2 Human-Centered Assessment of Text–Image Alignment
Table 2: Quantitative evaluation on T2I-CompBench++ across eight compositional categories using three different backbones. Results are reported for baseline models, SOTA methods, and our variants. CARINOX achieves the strongest overall alignment, consistently surpassing both optimization- and exploration-based baselines. Best values are in bold, and second-best are underlined.
5.3 Category-Level Compositional Benchmarking — T2I-CompBench++
We further benchmark our approach against a broad set of SOTA methods on T2I-CompBench++, which evaluates eight compositional categories using specialized evaluators for each dimension. Since different metrics are used across categories (e.g., CLIP for non-spatial relations), the absolute ranges vary, but the comparisons remain consistent across methods. Our baselines cover multiple methodological families: (i) multi-step open-source diffusion backbones (SD v1.4, SD v2.1, SDXL, PixArt-$\alpha$), (ii) single-step backbones (SD-Turbo, SDXL-Turbo, PixArt-$\alpha$ DMD), (iii) recent commercial systems such as DALL·E 3, (iv) attention-based inference-time control methods (Structured Diffusion Feng et al. (2023b), Attend-and-Excite Chefer et al. (2023)), (v) gradient-based initial noise optimization approaches (InitNO Guo et al. (2024b), ReNO Eyring et al. (2024)), and (vi) exploration strategies (Pick-a-Pic Kirstain et al. (2023), ImageSelect Karthik et al. (2023)). We report results for three variants of our method: CARINX (exploration only), CARINO (optimization only), and CARINOX (combined).
Importantly, CARINO and CARINX also provide consistent improvements when used independently: CARINX surpasses existing exploration methods, while CARINO establishes a new strong baseline for noise optimization. Together, the full CARINOX pipeline further amplifies these gains, outperforming commercial systems such as DALL·E 3 and advancing the SOTA across compositional categories.
@@ 6.2 Evaluation of Individual Reward Functions
Table 4: Effect of individual reward components on SD-Turbo evaluated on T2I-CompBench++. Single rewards might get category-specific gains, while the full CARINO combination achieves the best overall mean and the most balanced improvements across categories.
@@ C.1 Evaluation Metrics
DA Score (Singh and Zheng, 2023) 0.772 0.463 0.711 0.318 0.453 0.488 0.297 0.462
@@
DA Score (Singh and Zheng, 2023) 0.603 0.337 0.534 0.247 0.357 0.364 0.206 0.347
@@
DA Score (Singh and Zheng, 2023) 4
@@
VQA-based metrics assess compositional alignment by casting text–image consistency as a question answering task. Questions derived from the prompt are posed to a pretrained VQA model, with scores based on the correctness of its responses. VQAScore (Lin et al., 2024b) generates yes/no questions from the text, while TIFA (Hu et al., 2023) uses structured templates to cover objects, attributes, and relations. Variants target specific aspects: DA Score (Singh and Zheng, 2023) asks entity–attribute questions to test binding, DSG (Cho et al., 2023) converts the text into a scene graph to verify entities and relations, and B-VQA (Huang et al., 2023) decomposes the text into object–attribute pairs, querying each with BLIP-VQA and combining the probabilities.